dojaxlyah710:

given polygon abcd is similar to polygon qrst if ad = 17 and qt = 12. the perimeter of polygon abcd is 51, what is the perimeter of polygon qrst?

surjithayer:

polygons are similar so corresponding sides and perimeters are proportional. \[\frac{ qt }{ ad }=\frac{ P }{ 51 }\] perimeter of qrst P \[=\frac{ 12 }{ 17}\times 51=36 ~units\]